Luigu plans on investing 4000 coins he has collected from his travels into two accounts. The grumpy Goomba account earns 3% simple interest and the Bashful boo account earns 7% simple interest. How much money should he invest into the Bashful boo account so he can earn 186 coins at the end of the year?
Answer by 1335(25)

Let's say you invest "x" coins in the Bashful boo account. You now have 4000-x coins left to invest in the grumpy Goomba account.
In terms of a formula you get: ".07x + .03(4000-x)=186.
To simplify this, multiply by 100 to remove the decimals. You now have: "7x+3(4000-x)=18600".
Next, distribute the 3: "7x+12000-3x=18600".
Now simplify the problem by combining like terms. (the Xs): 7x-3x equals 4x.
Now re-write the equation: 4x+12000=18600 and subtract "12000" from both sides and get: 4x=6600
Divide both sides by 4 to get the answer x=1650 coins.
He should invest 1650 coins in the Bashful boo account in order to earn 186 coins at the end of the year.